it takes dave 12 hours to create a sculpture of a dolphin. it takes his buddy rowdy 9 hours to create the same sculpture. if they work together how long does it take dave and rowdy to create this amazing work of art?
Answer by Boreal(15235)

Dave does 1/12 of the sculpture in 1 hour and x/12 in x hours
Rowdy does x/9 in x hours
(x/12)+(x/9)=1 full sculpture
multiply by 36
3x+4x=36
7x=36
x=36/7 hours together, and this may be put in decimal form or hours/min form.